# EKF-Based Radar-Inertial Odometry with Online Temporal Calibration

### v1

@@ Abstract
Accurate time synchronization between heterogeneous sensors is crucial for ensuring robust state estimation in multi-sensor fusion systems. Sensor delays often cause discrepancies between the actual time when the event was captured and the time of sensor measurement, leading to temporal misalignment (time offset) between sensor measurement streams. In this paper, we propose an extended Kalman filter (EKF)-based radar-inertial odometry (RIO) framework that estimates the time offset online. The radar ego-velocity measurement model, estimated from a single radar scan, is formulated to include the time offset for the update. By leveraging temporal calibration, the proposed RIO enables accurate propagation and measurement updates based on a common time stream. Experiments on multiple datasets demonstrated the accurate time offset estimation of the proposed method and its impact on RIO performance, validating the importance of sensor time synchronization. Our implementation of the EKF-RIO with online temporal calibration is available at https://github.com/spearwin/EKF-RIO-TC.
@@ I INTRODUCTION
Accurate and robust state estimation is crucial for the successful execution of autonomous missions using mobile robots or vehicles. Global navigation satellite system (GNSS) provides reliable estimation in typical outdoor environments, but its reliability degrades or it is unavailable in obstructed areas such as urban canyons or indoor environments. In such cases, alternative state estimation methods such as simultaneous localization and mapping (SLAM) or odometry using exteroceptive sensors (e.g., camera, LiDAR, and radar) are crucial to maintain reliable autonomy.
@@
Over the past 20 years, real-time state estimation methods using light-based sensors such as a camera and a LiDAR have significantly improved accuracy and robustness in diverse environments [1, 2, 3, 4]. Vision-based methods show notable performance across a wide range of conditions, despite relying on small and lightweight sensors. However, their performance significantly decreases in environments with lighting changes or visually featureless surfaces [5]. In contrast, LiDAR-based methods are resilient to lighting conditions, and can accurately capture the detailed structure of the surrounding environment over long distances. However, they struggle in environments with self-similarity, such as long corridors or structureless areas like flat planes [6]. Moreover, light-based sensors face limitations when exposed to small particles, such as snow, fog, or dust, due to their short wavelength.
@@
Figure 1: Illustration of the temporal misalignment between IMU and radar measurement streams, along with the corresponding EKF’s propagation and measurement update steps. The time offset between the sensors is denoted as $t_{d}$. The time $t^{\prime}$ corresponds to the IMU measurement and $t$ to the radar measurement.
@@
Recently, a radar has gained attention as a promising solution to address these challenges [7]. In particular, a frequency-modulated continuous wave (FMCW) millimeter-wave 4D radar not only penetrates small particles effectively due to its relatively long wavelength but also measures relative speed (i.e., doppler velocity) to surrounding environments through frequency modulation. This radar typically operates at 5-20 Hz, providing raw signal data, 3D point clouds with spatial information and Doppler velocity for each point, enabling the estimation of ego-velocity from a single radar scan [8]. However, due to its relatively low sensor rate, it is necessary to predict the movement between consecutive radar scans. To address this, sensor fusion with an IMU, which operates at a higher sensor rate of over 100 Hz, can be utilized. The IMU complements the movement between consecutive radar scans, improving overall estimation accuracy and robustness. Therefore, there has been increasing interest and active research in radar-inertial odometry (RIO), which tightly couples the IMU with the radar.
@@
Accurate time synchronization between heterogeneous sensors is crucial for data fusion. All sensors inherently experience delays, leading to a discrepancy between the actual time when the event was captured and the time recorded as the sensor measurement, as illustrated in Fig. 1. While IMU systems typically have minimal latency, radar systems experience more significant delays due to factors such as signal processing steps, including fast Fourier transform, beamforming techniques, detection algorithms, and elevation angle estimation. Additionally, inherent hardware delays, including the analog-to-digital converter start-time and transmission delay, further contribute to the overall latency of radar systems [9]. Differences in sensor delays lead to temporal misalignments, which can pose significant challenges in multi-sensor systems. To address this issue, special hardware triggers or manufacturer software support may be required, but not all sensors provide such functions [10]. The time offset, which represents the difference in delays between IMU and radar systems, can reach up to hundreds of milliseconds, posing significant challenges for accurate data fusion in RIO. Compared with LiDAR-IMU and camera-IMU systems, our experimental results show that radar-IMU systems have a significantly larger time offset, highlighting the importance of temporal calibration.
@@
    The proposed method is validated through both self-collected dataset from real-world environments and open datasets, with and without hardware triggers. The results show that the time offset between sensors is non-negligible and must be accurately estimated to improve RIO performance; and
@@ II-A Radar-Inertial Odometry with FMCW 4D Radars
Doer and Trommer [11] proposed an EKF-based RIO framework that fuses IMU data with 3D ego-velocity estimated from radar measurements. They introduced the 3-point RANSAC-LSQ to remove outliers from the noisy radar measurements during the ego-velocity estimation. Later, the same authors extended the filter state to include the extrinsic parameters between the IMU and the radar for online extrinsic estimation [12], and enhanced yaw estimation for indoor environments using Manhattan world assumptions [13]. Michalczyk et al. [14] focused on the relatively accurate radar measurements, specifically distance and Doppler velocity, utilizing them for filter updates. In addition to ego-velocity updates, they constructed point-to-point residuals between consecutive scans and leveraged the most informative dimension, which is distance, for the measurement update. The same authors extended this approach by incorporating multi-pose into the framework and including long-term observed points in the filter state to further improve performance [15]. Zhuang et al. [16] adopted graduated non-convexity for ego-velocity estimation and applied distribution-to-multi-distribution constraints for current scan and submap matching when utilizing point residuals for sparse radar points. They also incorporated scancontext for place revisits and loop closures via pose graphs. Kubelka et al. [10] compared scan matching-based methods (e.g., ICP, APD-GICP, and NDT) and radar ego-velocity-based methods (e.g., IMU attitude with radar ego-velocity and EKF-RIO [11]) across two distinct radar sensor setups. Scan matching-based methods showed better performance in dense point cloud setups, while radar ego-velocity-based methods excelled in sparse point cloud setups. In sparse radar setups, scan matching-based methods experienced random divergence due to incorrect scan matching caused by the low radar point cloud density. Performance degradation in radar ego-velocity-based methods was also observed in specific sections, such as when the vehicle hit a bump, causing discrepancies between the IMU and the radar measurements due to the temporal misalignment. Their analysis on this issue motivated our research.
In [11, 12, 13], hardware triggers were implemented on a microcontroller board for sensor time synchronization. In [10], and [14, 15, 16], the authors did not account for the time offset when constructing their system frameworks, assuming the times of sensor measurements were accurate. Our proposed method estimates the time offset in real-time from radar ego-velocity, without relying on physical triggers. Since the ego-velocity is estimated from a single scan, there is no need for matching between consecutive scans, making the method independent of radar point cloud density, offering greater flexibility for various radar sensors. Furthermore, since all the mentioned works [10, 12, 11, 13, 14, 15, 16] utilize radar ego-velocity in the measurement update, the proposed method can be seamlessly integrated into their frameworks, not only ensuring accurate time synchronization but also potentially improving scan matching accuracy.
@@ II-B Online Temporal Calibration for Multi-Sensor Fusion Systems
Qin and Shen [17] proposed an optimization-based method for visual-inertial odometry (VIO) that enables online temporal calibration. They addressed time synchronization by jointly optimizing a prior factor, an IMU propagation factor, and a vision factor that accounts for the time offset between sensors. In particular, they compensated for feature measurements using the feature velocity on the image plane along with the time offset. Li and Mourikis [18] proposed a filter-based method to estimate the time offset between a camera and an IMU by including the time offset variable in the filter state. They formulated the camera measurement models for the 3D feature positions using the filter state including the time offset, thereby effectively estimating the time offset and improving the performance of VIO. Lee et al. [19] proposed a filter-based method for multi-sensor fusion odometry involving a LiDAR and an IMU. Since finding point correspondences between LiDAR scans can be challenging, they used plane patches, which contain structural information, to handle the data more efficiently. Their method involved extracting plane patches from the LiDAR point clouds and incorporating the time offset variable into the plane measurement models.
In contrast to previous works [17, 18, 19] which estimate the time offset by matching features between consecutive images or scans, our proposed method estimates the time offset from a single radar scan by formulating the radar ego-velocity measurement models. Unlike the camera and the LiDAR, the radar uniquely measures the Doppler velocity, enabling the direct estimation of ego-velocity. Instead of relying on scan matching, which can be challenging due to the sparse nature of radar point clouds, the proposed method avoids potential inaccuracies associated with correspondence matching.
@@ III Notation
Uppercase letters in superscripts (e.g., $A$ in ${}^{A}\mathbf{q}_{B}$) denote the reference coordinate frame. Quaternions, which are commonly used to represent attitude, follow the Hamilton convention [20]. Vectors are represented by bold lowercase letters, matrices by bold uppercase letters, and scalars by non-bold lowercase letters.
${}^{A}\mathbf{q}_{B}$ represents the quaternion describing the attitude of frame $B$ relative to frame $A$. The rotation matrix obtained from this quaternion, ${}^{A}\mathbf{R}_{B}=\mathbf{R}({}^{A}\mathbf{q}_{B})$, is the part of the special orthogonal group, $\text{SO}(3)$. ${}^{A}\mathbf{p}_{B}$ represents the position vector of frame $B$ relative to frame $A$, expressed in the $A$ frame.
@@ IV Filter Description
The system is represented with three coordinate frames: the global frame $G$, the IMU frame $I$, and the radar frame $R$. The proposed EKF-based RIO aims to estimate the 6D pose of the IMU-fixed frame $I$ with respect to the global reference frame $G$. The estimator utilizes the error state extended Kalman filter (ES-EKF), which is well-suited for handling the non-linear dynamics and measurement models typically encountered in pose estimation problems. By maintaining a minimal error-state and operating close to the origin, the ES-EKF avoids issues such as over-parameterization and singularities, ensuring consistency and efficiency. It also simplifies Jacobian computations, enhancing the filter’s robustness and computational efficiency when fusing IMU and radar measurements.
@@ IV-A System Overview
Figure 1 illustrates the temporal misalignment between IMU and radar streams, along with the corresponding execution of EKF’s propagation and update steps. The upper plot shows the actual time when the event was captured by the sensors, while the lower plot represents the recorded time of the sensor measurement. Each sensor measures an actual event at a certain time, but due to delays (i.e., $t_{d,IMU}$ and $t_{d,Radar}$), the sensor measurement reflects a later time. The time offset $t_{d}$ represents the difference between the delays of the IMU and the radar, defined as:
@@
Traditional EKF-based RIO performs propagation using IMU measurements until the radar measurement arrives, at which point the system executes the measurement update based on the times recorded in the sensor measurements. To ensure accurate state estimation, it is crucial to align the sensor measurements from both the IMU and the radar to a common time stream. While the exact delays of individual sensors are difficult to determine, the time offset $t_{d}$ can be estimated in real-time using the radar ego-velocity, allowing the system to adjust the radar measurement to align with the IMU measurement time stream, which serves as the common time reference. By leveraging temporal calibration, the proposed RIO enables propagation and measurement updates to be performed based on a common time stream.
The system state and its representation are explained in Section IV-B. In Section IV-C, we cover the propagation using the IMU, and in Section IV-D, the radar measurement update is discussed. The online temporal calibration is detailed in Section IV-E.
@@ IV-C Propagation with IMU Measurements
where ${}^{G}\mathbf{g}$ represents the gravity vector in the global frame. The estimated angular velocity ${}^{I}\hat{\bm{\omega}}_{I}$ and acceleration ${}^{I}\hat{\mathbf{a}}_{I}$ are expressed as ${}^{I}{\hat{\bm{\omega}}}_{I}={}^{I}{\bm{\omega}}_{I}^{m}-\hat{\mathbf{b}}_{g}$ and ${}^{I}\hat{\mathbf{a}}_{I}={}^{I}{\mathbf{a}}_{I}^{m}-\hat{\mathbf{b}}_{a}$, where ${}^{I}{\bm{\omega}}_{I}^{m}$ and ${}^{I}{\mathbf{a}}_{I}^{m}$ denote the gyroscope and accelerometer measurements, respectively, in the IMU frame. The matrix $\mathbf{\Omega}(\hat{\bm{\omega}})$, constructed from the estimated angular velocity $\hat{\bm{\omega}}$ and its skew-symmetric matrix $\lfloor\hat{\bm{\omega}}\times\rfloor$, is represented as:
@@
where $\mathbf{n}=\left(\mathbf{n}_{g}^{\top},\mathbf{n}_{wg}^{\top},\mathbf{n}_{a}^{\top},\mathbf{n}_{wa}^{\top},n_{d}\right)^{\top}$. The noise vectors $\mathbf{n}_{g}$ and $\mathbf{n}_{a}$ represent the Gaussian noise affecting the gyroscope and accelerometer measurements, respectively. Similarly, $\mathbf{n}_{wg}$ and $\mathbf{n}_{wa}$ correspond to the random walks for the gyroscope and accelerometer measurement biases. The term $n_{d}$ accounts for the Gaussian noise (i.e., uncertainty) in the time offset.
The matrix $\mathbf{F}$ represents the linearized system dynamics, and the matrix $\mathbf{G}$ models the influence of the process noise on the error state. Only the non-zero elements of the matrix $\mathbf{F}$ are given as follows:
@@
To propagate the covariance, the discrete-time state transition matrix $\mathbf{\Phi}_{k}$ and discrete-time process noise covariance matrix $\mathbf{Q}_{k}$, derived from Eq. (7), are defined as follows:

### v2

@@ Abstract
Accurate time synchronization between heterogeneous sensors is crucial for ensuring robust state estimation in multi-sensor fusion systems. Sensor delays often cause discrepancies between the actual time when the event was captured and the time of sensor measurement, leading to temporal misalignment (time offset) between sensor measurement streams. In this paper, we propose an extended Kalman filter (EKF)-based radar-inertial odometry (RIO) framework that estimates the time offset online. The radar ego-velocity measurement model, derived from a single radar scan, is formulated to incorporate the time offset into the update. By leveraging temporal calibration, the proposed RIO enables accurate propagation and measurement updates based on a common time stream. Experiments on both simulated and real-world datasets demonstrate the accurate time offset estimation of the proposed method and its impact on RIO performance, validating the importance of sensor time synchronization. Our implementation of the EKF-RIO with online temporal calibration is available at https://github.com/spearwin/EKF-RIO-TC.
@@ I INTRODUCTION
Accurate and robust state estimation is crucial for the successful execution of autonomous missions using mobile robots or vehicles. Global navigation satellite system provides reliable estimation in typical outdoor environments, but its reliability degrades or it is unavailable in obstructed areas such as urban canyons or indoor environments. In such cases, alternative state estimation methods such as simultaneous localization and mapping (SLAM) or odometry using exteroceptive sensors (e.g., camera, LiDAR, and radar) are crucial to maintain reliable autonomy.
@@
Real-time state estimation using light-based sensors such as cameras and LiDAR has significantly improved over the past two decades [1, 2, 3, 4]. Vision-based methods show notable performance across a wide range of conditions, despite their reliance on small and lightweight sensors. However, their performance significantly decreases in environments with lighting changes or visually featureless surfaces [5]. In contrast, LiDAR-based methods are resilient to lighting conditions and can accurately capture the detailed structure of the surrounding environment over long distances. However, they struggle in self-similar or structureless environments, such as long corridors or flat planes [6]. Moreover, light-based sensors face limitations when exposed to small particles, such as snow, fog, or dust, due to their short wavelength.
@@
Figure 1: Illustration of the temporal misalignment between IMU and radar measurement streams, along with the corresponding EKF propagation and measurement update steps. The time offset between sensors is denoted as $t_{d}=t_{d,IMU}-t_{d,Radar}$. For simplicity, the IMU delay is omitted in the figure. Two approaches are shown: with temporal calibration (proposed) and without it.
@@
Recently, a radar has gained attention as a promising solution to address these challenges [7]. In particular, a frequency-modulated continuous wave (FMCW) millimeter-wave 4D radar not only penetrates small particles effectively due to its relatively long wavelength but also measures relative speed (i.e., doppler velocity) to surrounding environments through frequency modulation. This radar typically provides measurements at a rate of 5-20 Hz, including raw signal data, 3D point clouds with spatial information and Doppler velocity for each point, enabling the estimation of ego-velocity from a single radar scan [8]. However, due to its relatively low sensor rate, it is necessary to predict the movement between consecutive radar scans. To address this, sensor fusion with an IMU, which operates at a higher sensor rate of over 100 Hz, can be utilized. The IMU complements the movement between consecutive radar scans, improving overall estimation accuracy and robustness. Therefore, there has been increasing interest and active research in radar-inertial odometry (RIO), including both loosely and tightly coupled methods, which integrate the IMU with the radar.
@@
Accurate time synchronization between heterogeneous sensors is crucial for data fusion. All sensors inherently experience delays, leading to a discrepancy between the actual time when the event was captured and the time when the measurement is timestamped, as illustrated in Fig. 1. While IMU systems typically have minimal latency, radar systems experience more significant delays due to factors such as signal processing steps, including fast Fourier transform, beamforming techniques, detection algorithms, and elevation angle estimation. Additionally, inherent hardware delays, including the analog-to-digital converter start-time and transmission delay, further contribute to the overall latency of radar systems [9]. Differences in sensor delays lead to temporal misalignments, which can pose significant challenges in multi-sensor systems. To address this issue, special hardware triggers or manufacturer software support may be required, but not all sensors provide such functions [10]. The time offset, which represents the difference in delays between IMU and radar systems, can reach up to hundreds of milliseconds, posing significant challenges for accurate data fusion in RIO. Compared with LiDAR-IMU and camera-IMU systems, our experimental results show that radar-IMU systems have a significantly larger time offset, highlighting the importance of temporal calibration.
@@
    The proposed method is validated through both simulated and real-world datasets with and without hardware triggers. The results show that the time offset between sensors is non-negligible and must be accurately estimated to improve RIO performance; and
@@ II-A Radar-Inertial Odometry with FMCW 4D Radars
Doer and Trommer [11] proposed an EKF-based RIO framework that fuses IMU data with 3D ego-velocity estimated from radar measurements. They introduced the 3-point RANSAC-LSQ to remove outliers in ego-velocity estimation. Later, the same authors extended the filter state to include the extrinsic parameters between the IMU and the radar for online extrinsic estimation [12], and enhanced yaw estimation for indoor environments using Manhattan world assumptions [13]. Michalczyk et al. [14] focused on the relatively accurate radar measurements, specifically distance and Doppler velocity, utilizing them for filter updates. In addition to ego-velocity updates, they constructed point-to-point residuals using distance as the most informative dimension for updates. The same authors extended this approach by incorporating multi-pose into the framework and including long-term observed points in the filter state to further improve performance [15]. Zhuang et al. [16] adopted graduated non-convexity for ego-velocity estimation and applied distribution-to-multi-distribution constraints for sparse scan matching. They also incorporated scancontext for place revisits and loop closures via pose graphs. Kubelka et al. [10] compared scan matching-based methods (e.g., ICP, APD-GICP, and NDT) and radar ego-velocity-based methods (e.g., IMU attitude with radar ego-velocity and EKF-RIO [11]) across two distinct radar sensor setups. The former showed better in dense point cloud setups, while the latter outperformed in sparse setups. In the sparse radar setups, scan matching often failed due to low point cloud density. Performance degradation in radar ego-velocity-based methods was also observed in specific sections, such as when the vehicle hit a bump, causing discrepancies between the IMU and the radar measurements due to the temporal misalignment. Their analysis on this issue motivated our research.
In [11, 12, 13], hardware triggers were implemented on a microcontroller board for sensor time synchronization. In [10], and [14, 15, 16], the authors did not account for the time offset when constructing their system frameworks, assuming the times of sensor measurements were accurate. Our proposed method estimates the time offset in real-time from radar ego-velocity, without relying on physical triggers. Since the ego-velocity is estimated from a single scan, there is no need for matching between consecutive scans, making the method independent of radar point cloud density and thereby offering greater flexibility for various radar sensors. Furthermore, since all the mentioned works [10, 12, 11, 13, 14, 15, 16] utilize radar ego-velocity in the measurement update, the proposed method can be seamlessly integrated into their frameworks, not only ensuring accurate time synchronization but also potentially improving scan matching accuracy.
@@ II-B Online Temporal Calibration for Multi-Sensor Fusion Systems
Qin and Shen [17] proposed an optimization-based method for visual-inertial odometry (VIO) that enables online temporal calibration. They addressed time synchronization by jointly optimizing prior, IMU, and vision factors which account for the time offset between sensors. In particular, they compensated for feature measurements using their velocities on the image plane along with the time offset. Li and Mourikis [18] proposed a filter-based method to estimate the time offset between a camera and an IMU by including the time offset variable in the filter state. They formulated the camera measurement models for the 3D feature positions using the filter state including the time offset, thereby effectively estimating the time offset and improving the performance of VIO. Lee et al. [19] proposed a filter-based method for multi-sensor fusion odometry involving a LiDAR and an IMU. Since finding point correspondences between LiDAR scans can be challenging, they used plane patches to handle the data more efficiently. Their method involved extracting plane patches from the point clouds and incorporating the time offset variable into the plane measurement models.
In contrast to previous works [17, 18, 19] which estimate the time offset by matching features between consecutive images or scans, our proposed method estimates the time offset from a single radar scan by formulating the radar ego-velocity measurement models. Unlike the camera and the LiDAR, the radar uniquely measures the Doppler velocity, enabling the direct estimation of ego-velocity. Instead of relying on scan matching, which can be challenging with sparse radar data, the proposed method avoids potential inaccuracies associated with correspondence matching.
@@ III Notation
Uppercase letters in superscripts (e.g., $A$ in ${}^{A}\mathbf{q}_{B}$) denote the reference frame. Quaternions follow the Hamilton convention [20]. Vectors are represented by bold lowercase letters, matrices by bold uppercase letters, and scalars by non-bold lowercase letters.
${}^{A}\mathbf{q}_{B}$ represents the quaternion describing the attitude of frame $B$ relative to $A$. The corresponding rotation matrix ${}^{A}\mathbf{R}_{B}=\mathbf{R}({}^{A}\mathbf{q}_{B})$ belongs to $\text{SO}(3)$. ${}^{A}\mathbf{p}_{B}$ represents the position of frame $B$ relative to $A$, expressed in $A$.
@@ IV Filter Description
The system uses three coordinate frames: global $G$, IMU $I$, and radar $R$. The proposed EKF-based RIO estimates the 6D pose of frame $I$ with respect to the global reference frame $G$. The estimator utilizes the error state extended Kalman filter (ES-EKF), which effectively handles nonlinear dynamics and measurement models in pose estimation. By maintaining a minimal error-state near the origin, the filter avoids over-parameterization and singularities, and simplifies Jacobian computations, enhancing consistency and efficiency when fusing IMU and radar measurements.
@@ IV-A System Overview
Figure 1 illustrates the temporal misalignment between the IMU and radar measurement streams, and the corresponding EKF propagation and update steps. The upper plot shows the actual time when the event was captured by the sensors, while the lower plot represents the time of measurement arrival, which corresponds to the timestamp assigned to the sensor data. Due to individual sensor delays (i.e., $t_{d,IMU}$ and $t_{d,Radar}$), the timestamp assigned to the measurement does not match the actual event time. The time offset $t_{d}$ represents the difference between the sensor delays, defined as:
@@
Traditional EKF-based RIO, which does not account for temporal calibration, performs propagation with IMU data and updates using radar measurements based on the time of measurement arrival. To ensure accurate state estimation, it is crucial to align the sensor measurements from both the IMU and the radar to a common time stream. While the exact delays of individual sensors are difficult to determine, the time offset can be estimated. The proposed method estimates the time offset $t_{d}$ in real-time using radar ego-velocity, and adjusts the radar measurement to align with the IMU measurement time stream, which serves as the common time reference. By leveraging temporal calibration, the proposed RIO enables propagation and measurement updates to be performed based on a common time stream.
@@ IV-C Propagation with IMU Measurements
where ${}^{G}\mathbf{g}$ represents the gravity vector. The estimated angular velocity ${}^{I}\hat{\bm{\omega}}_{I}$ and acceleration ${}^{I}\hat{\mathbf{a}}_{I}$ are expressed as ${}^{I}{\hat{\bm{\omega}}}_{I}={}^{I}{\bm{\omega}}_{I}^{m}-\hat{\mathbf{b}}_{g}$ and ${}^{I}\hat{\mathbf{a}}_{I}={}^{I}{\mathbf{a}}_{I}^{m}-\hat{\mathbf{b}}_{a}$, where ${}^{I}{\bm{\omega}}_{I}^{m}$ and ${}^{I}{\mathbf{a}}_{I}^{m}$ denote the gyroscope and accelerometer measurements, respectively. The matrix $\mathbf{\Omega}(\hat{\bm{\omega}})$, constructed from the estimated angular velocity $\hat{\bm{\omega}}$ and its skew-symmetric matrix $\lfloor\hat{\bm{\omega}}\times\rfloor$, is represented as:
@@
where $\mathbf{n}=\left(\mathbf{n}_{g}^{\top},\mathbf{n}_{wg}^{\top},\mathbf{n}_{a}^{\top},\mathbf{n}_{wa}^{\top},n_{d}\right)^{\top}$. The vectors $\mathbf{n}_{g}$ and $\mathbf{n}_{a}$ represent the Gaussian noise on the gyroscope and accelerometer, while $\mathbf{n}_{wg}$ and $\mathbf{n}_{wa}$ denote their bias random walks. The scalar $n_{d}$ models Gaussian noise (i.e., uncertainty) in the time offset.
The matrix $\mathbf{F}$ represents the linearized system dynamics, and $\mathbf{G}$ models the influence of the process noise on the error state. Only the non-zero elements of the matrix $\mathbf{F}$ are given as follows:
@@
To propagate the covariance, the discrete-time transition matrix $\mathbf{\Phi}_{k}$ and process noise covariance matrix $\mathbf{Q}_{k}$, derived from Eq. (7), are defined as follows: